### v1

# Linking Potentially Misclassified Healthy Food Access to Diabetes Prevalence

Abstract
Access to healthy food is key to maintaining a healthy lifestyle and can be quantified by the distance to the nearest grocery store. However, calculating this distance forces a trade-off between cost and correctness. Accurate route-based distances following passable roads are cost-prohibitive, while simple straight-line distances ignoring infrastructure and natural barriers are accessible yet error-prone. Categorizing low-access neighborhoods based on these straight-line distances induces misclassification and introduces bias into standard regression models estimating the relationship between disease prevalence and access. Yet, fully observing the more accurate, route-based food access measure is often impossible, which induces a missing data problem. We combat bias and address missingness with a new maximum likelihood estimator for Poisson regression with a binary, misclassified exposure (access to healthy food within some threshold), where the misclassification may depend on additional error-free covariates. In simulations, we show the consequence of ignoring the misclassification (bias) and how the proposed estimator corrects for bias while preserving more statistical efficiency than the complete case analysis (i.e., deleting observations with missing data). Finally, we apply our estimator to model the relationship between census tract diabetes prevalence and access to healthy food in northwestern North Carolina.
Keywords: Food Environment, Maximum Likelihood Estimation, Two-Phase Design, Missing Data, One-Sided Misclassification, Poisson Regression
1 Introduction
Maintaining a healthy diet is more than simply a choice. It requires a person to have consistent access to and ability to obtain healthy food. Physical access to healthy foods may be hindered by many factors, including the neighborhood’s geography or community’s socioeconomic makeup (Wolfson et al., 2019). Subsequently, neighborhoods with poor food access may see higher prevalences of diseases linked to healthy eating (Gucciardi et al., 2014). We focus on the relationship between diabetes prevalence and food access at the neighborhood (census tract) level, although our methods adapt to other outcomes as well.
To estimate this relationship, we must first quantify food access, for which there are multiple metrics. We devote our attention to “access indicators,” which classify neighborhoods into binary categories of having/not having access to healthy food within some fixed distance. The United States Department of Agriculture (USDA) includes indicators like these in their Food Research Atlas (U.S. Department of Agriculture (USDA), 2019). Computing these food access indicators requires researchers to first calculate the pairwise distances between neighborhoods and potential healthy food retailers. Just like there are multiple ways to measure food access, there are multiple ways to measure these distances. For instance, the Haversine formula modifies the traditional straight-line Euclidean distance between two points (e.g., a neighborhood centroid and grocery store) to account for the curvature of the Earth. Alternatively, route-based distance calculations are possible, which only allow travel via established infrastructure (like roads) to connect the two points.
These methods each have their advantages, and choosing between them to calculate distances requires a trade-off. The Haversine formula is a simple mathematical equation, making it free and easy to calculate in existing software (e.g., Hijmans, 2022). However, it underestimates distances from neighborhoods to healthy food retailers and therefore overestimates neighborhoods’ food access. This underestimation stems from the fact that straight-line distances are necessarily less than or equal to any other measure of distance. In our study, underestimated Haversine distances can lead to neighborhoods being misclassified as nominally having access to healthy foods when they truly do not (false positives). Specifically, the access indicators are subject to one-sided misclassification, since false positives are the only possible type of error.
In short, Haversine distance is easily obtained but leads to access indicators that may be incorrect. Taking route-based distances via software like Google Maps circumvents the possibility of unrealistic, impassable routes. Still, these more-accurate distances are computationally expensive and, despite availability of open source options, may also be financially expensive. Therefore, calculating them for all neighborhoods may not be plausible.
Luckily, some accurate data are better than none, so we frame our study as a two-phase design. We measure the diabetes prevalence and the error-prone food access indicator based on Haversine distances for all neighborhoods (Phase I). Then, we query a subset of neighborhoods using Google Maps (Kahle and Wickham, 2013) to obtain the route-based distances for the more-accurate food access indicator (Phase II). We leverage missing data techniques to include every neighborhood in the model in some way, and this strategy yields unbiased estimates of the exposure effects (food access) on the outcome (diabetes prevalence) without fully observing the route-based exposures. For extensive reviews of statistical methods for error-prone data, see Keogh et al. (2020) and Shaw et al. (2020).
For our study, with the prevalence of diabetes as the outcome and the indicator of food access as the misclassified exposure, a new maximum likelihood estimator (MLE) for Poisson regression with exposure misclassification is needed. The rest of this paper discusses the derivation of the MLE (Section 2), extensive simulations (Section 3), and a case study (Section 4). We conclude by discussing our findings and next steps (Section 5).
2 Methods
2.1 Model and Notation
We are interested in modeling $Y$, a count outcome offset by $O$, given a binary exposure $X$ and covariate vector $\boldsymbol{Z}$. An error-prone, potentially misclassified version of the exposure, $X^{*}$, is available for all $N$ neighborhoods in the sample, while true $X$ is only available for a subset of $n$ “queried” neighborhoods ($n<N$). We can write the joint probability of a complete (queried) observation as follows:
${\Pr}_{\boldsymbol{\beta},\boldsymbol{\eta}}(Y,X,X^{*},\boldsymbol{Z})$ $={\Pr}(Y\mid X,X^{*},\boldsymbol{Z}){\Pr}_{\boldsymbol{\eta}}(X\mid X^{*},\boldsymbol{Z}){\Pr}(X^{*},\boldsymbol{Z})$
$={\Pr}_{\boldsymbol{\beta}}(Y\mid X,\boldsymbol{Z}){\Pr}_{\boldsymbol{\eta}}(X\mid X^{*},\boldsymbol{Z}){\Pr}(X^{*},\boldsymbol{Z}).$ (1)
First, ${\Pr}_{\boldsymbol{\beta}}(Y\mid X,\boldsymbol{Z})$ denotes the Poisson regression for $Y$ given ($X$,$\boldsymbol{Z}$), parameterized by $\boldsymbol{\beta}$. Second, ${\Pr}_{\boldsymbol{\eta}}(X\mid X^{*},\boldsymbol{Z})$ denotes the conditional probability mass function (PMF) of the true $X$ given error-prone $X^{*}$ and additional covariates $\boldsymbol{Z}$. Since $X$ is binary, we estimate this misclassification mechanism using logistic regression, parameterized by $\boldsymbol{\eta}$. To accommodate one-sided misclassification (e.g., only false positives), this model must be modified (Web Appendix A.1) but the rest of the methods remain unchanged. Third, ${\Pr}(X^{*},\boldsymbol{Z})$ denotes the joint marginal probability density or mass function of the error-prone and additional covariates $(X^{*},\boldsymbol{Z})$. Since $(X^{*},\boldsymbol{Z})$ are fully observed, ${\Pr}(X^{*},\boldsymbol{Z})$ drops out of our log-likelihood for $\boldsymbol{\beta}$, so we can leave this model unspecified. Notice that the second equality in (1) assumes conditional independence of ($Y$,$X^{*}$) given ($X$,$\boldsymbol{Z}$) (i.e., surrogacy).
In our setting, all $N$ neighborhoods have an observed $X^{*}$ based on Haversine distances, but only a limited subset of size $n$ have a measurement for $X$ based on true route-based distances. Thus, the data available for a neighborhood depend on whether it was queried:
Observed data =$=\begin{cases}(Y,X,X^{*},\boldsymbol{Z})&\text{ if the neighborhood was queried, and }\\ (Y,X^{*},\boldsymbol{Z})&\text{ otherwise.}\end{cases}$
Queried neighborhoods are assumed to follow (1), while unqueried ones are assumed to follow ${\Pr}_{\boldsymbol{\beta},\boldsymbol{\eta}}(Y,X^{*},\boldsymbol{Z})=\sum\limits_{x=0}^{1}{\Pr}_{\boldsymbol{\beta}}(Y\mid X=x,\boldsymbol{Z}){\Pr}_{\boldsymbol{\eta}}(X=x\mid X^{*},\boldsymbol{Z}){\Pr}(X^{*},\boldsymbol{Z})$ instead.
2.2 Maximum Likelihood Estimation
We piece the queried and unqueried observations and their joint probability functions together, along with an indicator $Q_{i}$ of queried status, to define the likelihood function:
$\mathcal{L}_{N}(\boldsymbol{\beta},\boldsymbol{\eta})$ $=\prod_{i=1}^{N}\{{\Pr}_{\boldsymbol{\beta},\boldsymbol{\eta}}(X_{i},X_{i}^{*},Y_{i},\boldsymbol{Z}_{i})\}^{Q_{i}}\{{\Pr}_{\boldsymbol{\beta},\boldsymbol{\eta}}(X_{i}^{*},Y_{i},\boldsymbol{Z}_{i})\}^{1-Q_{i}}$
$\propto\prod_{i=1}^{N}\{{\Pr}_{\boldsymbol{\beta}}(Y_{i}\mid X_{i},\boldsymbol{Z}_{i}){\Pr}_{\boldsymbol{\eta}}(X_{i}\mid X_{i}^{*},\boldsymbol{Z}_{i})\}^{Q_{i}}$
×$\times\left\{\sum_{x=0}^{1}{\Pr}_{\boldsymbol{\beta}}(Y_{i}\mid X_{i}=x,\boldsymbol{Z}_{i}){\Pr}_{\boldsymbol{\eta}}(X_{i}=x\mid X_{i}^{*},\boldsymbol{Z}_{i})\right\}^{1-Q_{i}}.$ (2)
The query indicator controls which probability function governs observation $i$ in (2); it evaluates to $1$ for queried neighborhoods (i.e., when $X_{i}$ is available) and to 00 otherwise. Notice that the $N$ neighborhoods are assumed to be independent in this likelihood. Modifying the likelihood to allow spatial autocorrelation is a promising future direction.
The MLEs $\hat{\boldsymbol{\beta}}$ maximize (2) and are found via an EM algorithm (Web Appendix A.2). Standard errors (SEs) for $\hat{\boldsymbol{\beta}}$ are conveniently estimated using numerical derivatives. Our proposed MLEs are implemented in the possum R package, available at www.github.com/sarahlotspeich/possum. Under regularity conditions, missingness at random, and proper specification of both models, the MLEs are asymptotically consistent, normal, and efficient.
3 Simulations
We conducted simulation studies to test the performance of our MLE against the gold standard, naive, and complete case analyses. The gold standard analysis fit a Poisson regression using $X$ for all observations, while the naive analysis used error-prone version $X^{*}$, and the complete case used $X$ for only the queried subset. Data generation was designed to mimic our data from the Piedmont Triad (Section 4), with primary variations of the sample size $N$, query proportion $q$, and positive predictive value $PPV=\Pr(X=1|X^{*}=1,Z)$.
[…]
We focus here on simulating varied misclassification rates (Section 3.1) and query percentages (Section 3.2). Unless otherwise stated, we fixed $q=0.1$, $PPV=0.6$, $\beta_{1}=0.18$, and outcome prevalence $\approx 10\%$. Additional results with two-sided misclassification in $X^{*}$ and varied outcome prevalence can be found in Web Appendix B.2–B.3. Like in the cases below, the MLE remained unbiased and more efficient than the complete case analysis.
3.1 Results Under Varied Misclassification Rates
We considered $50\%$ to $90\%$ $PPV$ at sample sizes of $N=390$ and $2200$. The bias of the MLE remained comparable to that of the gold standard and complete case analyses in all settings (Table 1). Even under the most severe misclassification ($PPV=0.5$), the MLE was less than $0.1\%$ biased. However, the naive analysis yielded heavy downward bias ranging from $-8.9\%$ to $-46.5\%$, worsening as $PPV$ decreased and persisting in larger samples.
Table 1: Simulation results for $\hat{\beta}_{1}$, the adjusted log prevalence ratio on $X$, under varied positive predictive value ($PPV$) and sample size ($N$), assuming $\approx 10\%$ outcome prevalence, $\beta_{1}=0.18$, and $q=0.1$ query percentage.
Gold Standard Naive Complete Case Maximum Likelihood Estimator
$\boldsymbol{N}$ $\boldsymbol{PPV}$ Bias ESE Bias ESE Bias ESE RE Bias ESE ASE CP RE
390 0.5 $0.000$ $0.005$ $-0.465$ $0.008$ $0.002$ $0.015$ $0.098$ $0.000$ $0.005$ $0.005$ $0.949$ $0.955$
0.6 $0.000$ $0.005$ $-0.367$ $0.007$ $0.000$ $0.016$ $0.085$ $0.000$ $0.005$ $0.005$ $0.954$ $0.975$
0.7 $0.001$ $0.005$ $-0.271$ $0.007$ $-0.001$ $0.018$ $0.075$ $0.001$ $0.005$ $0.005$ $0.938$ $0.985$
0.8 $0.000$ $0.005$ $-0.181$ $0.007$ $0.000$ $0.019$ $0.059$ $0.000$ $0.005$ $0.005$ $0.953$ $0.985$
0.9 $0.000$ $0.005$ $-0.089$ $0.006$ $0.000$ $0.030$ $0.028$ $0.000$ $0.005$ $0.005$ $0.944$ $0.987$
2200 0.5 $0.000$ $0.002$ $-0.465$ $0.003$ $-0.001$ $0.006$ $0.107$ $0.000$ $0.002$ $0.002$ $0.943$ $0.986$
0.6 $0.001$ $0.002$ $-0.366$ $0.003$ $0.001$ $0.007$ $0.097$ $0.001$ $0.002$ $0.002$ $0.935$ $0.984$
0.7 $0.000$ $0.002$ $-0.272$ $0.003$ $-0.002$ $0.007$ $0.087$ $0.000$ $0.002$ $0.002$ $0.951$ $0.992$
0.8 $0.000$ $0.002$ $-0.179$ $0.003$ $0.000$ $0.008$ $0.065$ $0.000$ $0.002$ $0.002$ $0.948$ $1.002$
0.9 $-0.001$ $0.002$ $-0.089$ $0.003$ $-0.002$ $0.011$ $0.036$ $-0.001$ $0.002$ $0.002$ $0.944$ $0.992$
[…]
The empirical SEs of the MLE were smaller than those of the complete case analysis in every setting. In addition, the average SEs of the MLE approximated the empirical versions very well, which also led to empirical CP in the correct neighborhood of 95%. Finally, the relative efficiency of the MLE to the gold standard analysis was over five times higher than that of the complete case analysis in every setting (between 0.960.960.960.961.001.001.001.00 versus 0.040.040.040.040.110.110.110.11).
3.2 Results Under Varied Query Percentages
We also considered querying proportions between $q=10\%$ and $75\%$ from samples $N=390$ and $2200$ neighborhoods. The bias of the MLE remained comparable to that of the gold standard analysis; both methods were under 0.5% biased in all settings (Table 2). The empirical SEs of the MLE continued to match or beat those from the complete case analysis in every setting. With the smallest $q$, the SEs from the complete case analysis were over three times the size of the MLE’s. Again, the average SEs of the MLE matched their empirical counterparts, and empirical CP was close to the nominal 95% level.
Table 2: Simulation results for $\beta_{1}$ under varied query percentage ($q$) and sample size ($N$), assuming $\approx 10\%$ outcome prevalence, $\beta_{1}=0.18$, and positive predictive value $=0.6$.
Gold Standard Naive Complete Case Maximum Likelihood Estimator
$\boldsymbol{N}$ $\boldsymbol{q}$ Bias ESE Bias ESE Bias ESE RE Bias ESE ASE CP RE
390 0.10 $0.000$ $0.005$ $-0.367$ $0.007$ $0.000$ $0.016$ $0.085$ $0.000$ $0.005$ $0.005$ $0.954$ $0.975$
0.25 $0.002$ $0.005$ $-0.367$ $0.007$ $0.003$ $0.010$ $0.227$ $0.002$ $0.005$ $0.005$ $0.953$ $0.974$
0.50 $0.001$ $0.005$ $-0.366$ $0.007$ $0.002$ $0.007$ $0.463$ $0.001$ $0.005$ $0.005$ $0.953$ $0.992$
0.75 $0.000$ $0.005$ $-0.367$ $0.007$ $0.000$ $0.006$ $0.673$ $0.000$ $0.005$ $0.005$ $0.953$ $1.000$
2200 0.10 $0.001$ $0.002$ $-0.366$ $0.003$ $0.001$ $0.007$ $0.097$ $0.001$ $0.002$ $0.002$ $0.935$ $0.984$
0.25 $0.000$ $0.002$ $-0.368$ $0.003$ $0.000$ $0.004$ $0.227$ $0.000$ $0.002$ $0.002$ $0.949$ $0.980$
0.50 $0.000$ $0.002$ $-0.367$ $0.003$ $0.000$ $0.003$ $0.508$ $0.000$ $0.002$ $0.002$ $0.945$ $0.992$
0.75 $-0.001$ $0.002$ $-0.368$ $0.003$ $-0.001$ $0.002$ $0.695$ $-0.001$ $0.002$ $0.002$ $0.948$ $1.000$
[…]
The relative efficiency of the MLE to the gold standard analysis exceeded that of the complete case analysis in every setting. This trend was especially noticeable (as expected) when $q$ was smaller, when larger percentages of observations were missing error-free exposures. The complete case analysis would have thrown these observations away.
4 Case Study: Piedmont Triad, North Carolina
4.1 Data and Objectives
We investigated the associations between having access to healthy foods and diabetes prevalence across $N=387$ census tracts in the Piedmont Triad region of northwestern North Carolina. A census tract is a small, relatively permanent statistical subdivision of a county or statistically equivalent entity (United States Census Bureau, 2022). Approximately 1.7 million people live in the Piedmont Triad, and the region is known to have excellent transportation infrastructure. The model of interest is defined as $\log\left\{\textrm{E}_{\boldsymbol{\beta}}(Y\mid X,Z)\right\}=\beta_{0}+\beta_{1}X+\beta_{2}Z+\beta_{3}X\times Z+\log(O)$, where $Y$ is the number of diabetes cases, $O$ is the population (used as an offset), $X$ is an indicator of having access to healthy foods within a given radius, and $Z$ is an indicator of metropolitan status. We calculated the Haversine and route-based distances from the population-weighted centroid, according to the 2020 U.S. Census, to the nearest healthy food retailer. Using one-mile and half-mile radii, we discretized these distances to create our exposures $X$ and $X^{*}$. For example, $X=1$ if a census tract has at least one healthy food retailer within the threshold and $=0$ otherwise.
Queried data (i.e., with both $X$ and $X^{*}$) were available for all census tracts. However, to demonstrate the two-phase design, we assumed that only a subset had available route-based access data for the MLE and complete case analyses. After running the gold standard analysis on the full data, we created a partially queried version wherein $X$ was only available for census tracts with $X^{*}=0$ (such that $X=0$ necessarily) and a subset of those with $X^{*}=1$. A random sample of $n=77$ tracts with $X^{*}=1$, targeting a 50/50 split between metropolitan/non-metropolitan, was chosen. Only tracts with error-prone access (i.e., $X^{*}=1$) needed to be validated, since those with $X^{*}=0$ were already known to have $X=0$ and do not inform the misclassification mechanism. The naive analysis used the error-prone access variable on the entire sample, and the complete case analysis used error-free access variable on the subset of queried data. The MLE used the error-prone and error-free access indicators when available and just the error-prone version otherwise.
4.2 Diabetes Prevalence and Healthy Food Access
At both half- and one-mile radii, the majority of tracts in the Piedmont Triad ($n=353$ [$91\%$] and $n=265$ [$68\%$], respectively) did not have access to healthy foods (Figure 1). Larger, rural tracts tended not to have access. Smaller, urban tracts tended to have access, or at least to be misclassified as such based on Haversine distances. Of the $192$ tracts with access to healthy foods based on one-mile Haversine distances, only $122$ actually had it based on route-based ones ($PPV=122/192=64\%$). As expected, there was more misclassification at the smaller threshold ($PPV=34/85=40\%$).
Figure 1: Map of Piedmont Triad census tracts’ access at one-half and one mile thresholds.
The estimated prevalence ratio $\widehat{PR}_{non}=\exp(\hat{\beta_{1}})$ compared the predicted prevalence of diabetes in non-metropolitan census tracts with versus without access to healthy foods. In metropolitan census tracts, $\widehat{PR}_{metro}=\exp(\hat{\beta_{1}}+\hat{\beta_{3}})$ was interpreted similarly, and 95% CIs for this prevalence ratio were calculated as $\exp\left\{(\hat{\beta_{1}}+\hat{\beta_{3}})\pm 1.96\widehat{\textrm{SE}}(\hat{\beta_{1}}+\hat{\beta_{3}})\right\}$, where $\widehat{\textrm{SE}}(\hat{\beta_{1}}+\hat{\beta_{3}})=\sqrt{\widehat{\textrm{V}}(\hat{\beta_{1}})+\widehat{\textrm{V}}(\hat{\beta_{3}})+\widehat{\textrm{Cov}}(\hat{\beta_{1}},\hat{\beta_{3}})}$. We focus on these two prevalence ratios; the full models are summarized in Supplemental Table S4.
At either threshold, prevalence of diabetes in neighborhoods with access to healthy foods was expected to be higher than in those without (Figure 2). This counterintuitive trend held at both distance thresholds and across all analysis methods. For example, based on the gold standard, diabetes prevalence in metropolitan tracts with access to healthy foods at one mile was expected to be 18% higher than in those without ($\widehat{PR}_{metro}=1.18$; 95% CI: $1.16$, $1.21$); among non-metropolitan tracts, those with access were expected to have 9% higher prevalence ($\widehat{PR}_{non}=1.09$; 95% CI: $1.06$, $1.11$).
Figure 2: Prevalence ratio estimates and 95% confidence intervals for healthy food access at a one-half and one mile threshold in the Piedmont Triad using four analysis methods.
The difference between tracts with and without access was more pronounced in metropolitan areas, according to the gold standard analyses. At both thresholds, the MLE also captured this difference, while the complete case analysis only did so at the one-half mile radius, and the naive analysis did not catch it at either threshold. The MLE actually exaggerated this difference at the one-mile threshold but otherwise closely aligned with the gold standard. Its 95% confidence intervals were slightly narrower than the complete case. For more details, see Web Appendix C.
5 Conclusion
We introduced a novel maximum likelihood estimator for Poisson regression with a misclassified binary exposure. In simulations, we observe that the MLE, on average, corrects for bias induced by the misclassified exposure in the naive analysis and reduces variance induced by deleting unqueried data in the complete case analysis. In our analysis of the Piedmont Triad, tracts with access to healthy foods had higher diabetes prevalence, on average, than those without. This difference was more pronounced for metropolitan tracts.
Our approach maximizes available information and addresses some of the concerns that arise in competing approaches. In simulations, the MLE had lower bias than the naive analysis, and it recovered much of the efficiency lost by the complete case analysis. For broader applicability, our MLE handles both one- and two-sided exposure misclassification.
We are potentially limited by our choice to assume a parametric logistic regression model for the exposure error mechanism. However, this choice has been shown to be fairly robust to misspecification (Lotspeich et al., 2022). Still, nonparametric or semiparametric error models are promising. Results from the case study are limited by the nature of the available data. Census tracts do not always fairly describe neighborhoods, but they were the most granular level of available diabetes data. Additionally, the model used to generate the small area diabetes estimates incorporates many socioeconomic factors that may have been relevant to our analysis but could not be included because of “double dipping.”
There are many interesting avenues of further exploration. It may be informative to explore spatial relationships among adjacent tracts, vary the parametric form of the outcome model, or modify the MLE to handle other types of data. The MLE’s efficiency could be further improved by a more strategic query design. Targeted strategies for selecting the most informative observations are common in the two-phase design literature.
SUPPLEMENTARY MATERIAL
Supporting information:
Web Appendices and Supplemental Figures and Tables (.pdf)
Data:
R package:
R package possum implementing the MLE (GNU zipped tar file)
R code:
R scripts for the simulation studies and Piedmont Triad analysis (.zip file)
Additional materials:
Additional code and results available for reproducibility at
https://github.com/ashleymullan/food_access_misclassification.

### v2

@@
 Maintaining a healthy diet is more than simply a choice.
-It requires a person to have consistent access to and ability to obtain healthy food.
+A person must have consistent access to and ability to obtain healthy food.
 Physical access to healthy foods may be hindered by many factors, including the neighborhood’s geography or community’s socioeconomic makeup (Wolfson et al., 2019).
-Subsequently, neighborhoods with poor food access may see higher prevalences of diseases linked to healthy eating (Gucciardi et al., 2014).
-We focus on the relationship between diabetes prevalence and food access at the neighborhood (census tract) level, although our methods adapt to other outcomes as well.
+Subsequently, neighborhoods with poor food access may see higher prevalences of diet-related diseases (Gucciardi et al., 2014).
+For example, Lopez (2007) links food access, specifically the presence of at least one grocery store within a zip code, to a 10% decrease in obesity risk when compared to zip codes without this access.
+Groups such as food policy councils (Lange et al., 2021) prioritize health equity by supporting policies that increase access to healthy food in rural and minority communities.
+We focus on the relationship between diabetes prevalence and food access at the neighborhood (census tract) level, although our methods adapt to other outcomes and units as well.
 To estimate this relationship, we must first quantify food access, for which there are multiple metrics.
@@
 The United States Department of Agriculture (USDA) includes indicators like these in their Food Research Atlas (U.S.
-Department of Agriculture (USDA), 2019).
+Department of Agriculture, 2019).
 Computing these food access indicators requires researchers to first calculate the pairwise distances between neighborhoods and potential healthy food retailers.
@@
 These methods each have their advantages, and choosing between them to calculate distances requires a trade-off.
-The Haversine formula is a simple mathematical equation, making it free and easy to calculate in existing software (e.g., Hijmans, 2022).
-However, it underestimates distances from neighborhoods to healthy food retailers and therefore overestimates neighborhoods’ food access.
+The Haversine formula is a simple mathematical equation, making it free, easy, and quick to calculate in existing software (e.g., Hijmans, 2022).
+However, it underestimates distances to healthy food retailers and therefore overestimates neighborhoods’ food access.
 This underestimation stems from the fact that straight-line distances are necessarily less than or equal to any other measure of distance.
-In our study, underestimated Haversine distances can lead to neighborhoods being misclassified as nominally having access to healthy foods when they truly do not (false positives).
-Specifically, the access indicators are subject to one-sided misclassification, since false positives are the only possible type of error.
+In our study, underestimated Haversine distances can lead to neighborhoods being misclassified as having access to healthy foods when they do not (false positives).
+Specifically, the access indicators are subject to one-sided misclassification; false positives are the only possible type of error.
 In short, Haversine distance is easily obtained but leads to access indicators that may be incorrect.
 Taking route-based distances via software like Google Maps circumvents the possibility of unrealistic, impassable routes.
-Still, these more-accurate distances are computationally expensive and, despite availability of open source options, may also be financially expensive.
+Still, these more-accurate distances are computationally expensive and, despite availability of open source options, may also be costly.
 Therefore, calculating them for all neighborhoods may not be plausible.
@@
 (2020).
-For our study, with the prevalence of diabetes as the outcome and the indicator of food access as the misclassified exposure, a new maximum likelihood estimator (MLE) for Poisson regression with exposure misclassification is needed.
+For our study, with diabetes prevalence as the outcome and (potentially mismeasured) food access indicator as the exposure, a new maximum likelihood estimator (MLE) for Poisson regression with exposure misclassification is needed.
+This MLE incorporates all information for all neighborhoods, offering better statistical efficiency than design-based approaches, like inverse probability weighting.
+As a model-based approach, it offers flexibility in choosing the queried subset; design-based approaches require nonzero sampling probability for all observations.
+Moreover, it is computationally faster than multi-stage model-based approaches, like multiple imputation or Bayesian modeling, and it would be straightforward to adapt for more robustness (e.g., semiparametric MLE).
 The rest of this paper discusses the derivation of the MLE (Section 2), extensive simulations (Section 3), and a case study (Section 4).
-We conclude by discussing our findings and next steps (Section 5).
+We conclude by discussing findings and next steps (Section 5).
 2 Methods
 2.1 Model and Notation
-We are interested in modeling $Y$, a count outcome offset by $O$, given a binary exposure $X$ and covariate vector $\boldsymbol{Z}$.
-An error-prone, potentially misclassified version of the exposure, $X^{*}$, is available for all $N$ neighborhoods in the sample, while true $X$ is only available for a subset of $n$ “queried” neighborhoods ($n<N$).
-We can write the joint probability of a complete (queried) observation as follows:
+We model $Y$, a count outcome offset by $O$, given a binary exposure $X$ and covariate vector $\boldsymbol{Z}$.
+An error-prone, potentially misclassified version of the exposure, $X^{*}$, is available for all $N$ sampled neighborhoods, while true $X$ is only available for a subset of $n$ “queried” ones ($n<N$).
+We write the joint probability of a complete (queried) observation as follows:
 ${\Pr}_{\boldsymbol{\beta},\boldsymbol{\eta}}(Y,X,X^{*},\boldsymbol{Z})$ $={\Pr}(Y\mid X,X^{*},\boldsymbol{Z}){\Pr}_{\boldsymbol{\eta}}(X\mid X^{*},\boldsymbol{Z}){\Pr}(X^{*},\boldsymbol{Z})$
 $={\Pr}_{\boldsymbol{\beta}}(Y\mid X,\boldsymbol{Z}){\Pr}_{\boldsymbol{\eta}}(X\mid X^{*},\boldsymbol{Z}){\Pr}(X^{*},\boldsymbol{Z}).$ (1)
-First, ${\Pr}_{\boldsymbol{\beta}}(Y\mid X,\boldsymbol{Z})$ denotes the Poisson regression for $Y$ given ($X$,$\boldsymbol{Z}$), parameterized by $\boldsymbol{\beta}$.
-Second, ${\Pr}_{\boldsymbol{\eta}}(X\mid X^{*},\boldsymbol{Z})$ denotes the conditional probability mass function (PMF) of the true $X$ given error-prone $X^{*}$ and additional covariates $\boldsymbol{Z}$.
-Since $X$ is binary, we estimate this misclassification mechanism using logistic regression, parameterized by $\boldsymbol{\eta}$.
+First, ${\Pr}_{\boldsymbol{\beta}}(Y\mid X,\boldsymbol{Z})$ denotes the Poisson regression for $Y$ given ($X$, $\boldsymbol{Z}$), parameterized by log prevalence ratios (log PRs) $\boldsymbol{\beta}$.
+Second, ${\Pr}_{\boldsymbol{\eta}}(X\mid X^{*},\boldsymbol{Z})$ denotes the conditional probability mass function for $X$ given ($X^{*}$, $\boldsymbol{Z}$).
+We estimate this misclassification mechanism using logistic regression, governed by $\boldsymbol{\eta}$.
 To accommodate one-sided misclassification (e.g., only false positives), this model must be modified (Web Appendix A.1) but the rest of the methods remain unchanged.
-Third, ${\Pr}(X^{*},\boldsymbol{Z})$ denotes the joint marginal probability density or mass function of the error-prone and additional covariates $(X^{*},\boldsymbol{Z})$.
-Since $(X^{*},\boldsymbol{Z})$ are fully observed, ${\Pr}(X^{*},\boldsymbol{Z})$ drops out of our log-likelihood for $\boldsymbol{\beta}$, so we can leave this model unspecified.
+Third, ${\Pr}(X^{*},\boldsymbol{Z})$ denotes the joint probability density or mass function of ($X^{*}$$\boldsymbol{Z}$).
+Since ($X^{*}$$\boldsymbol{Z}$) are fully observed, ${\Pr}(X^{*},\boldsymbol{Z})$ drops out of our log-likelihood for $\boldsymbol{\beta}$, so we can leave this model unspecified.
 Notice that the second equality in (1) assumes conditional independence of ($Y$, $X^{*}$) given ($X$, $\boldsymbol{Z}$) (i.e., surrogacy).
-In our setting, all $N$ neighborhoods have an observed $X^{*}$ based on Haversine distances, but only a limited subset of size $n$ have a measurement for $X$ based on true route-based distances.
+In our setting, all $N$ neighborhoods have an observed $X^{*}$ based on Haversine distances, but only a limited subset of size $n$ has true $X$ measured based on route-based distances.
 Thus, the data available for a neighborhood depend on whether it was queried:
-Observed data =$=\begin{cases}(Y,X,X^{*},\boldsymbol{Z})&\text{ if the neighborhood was queried, and }\\ (Y,X^{*},\boldsymbol{Z})&\text{ otherwise.}\end{cases}$
-Queried neighborhoods are assumed to follow (1), while unqueried ones are assumed to follow ${\Pr}_{\boldsymbol{\beta},\boldsymbol{\eta}}(Y,X^{*},\boldsymbol{Z})=\sum\limits_{x=0}^{1}{\Pr}_{\boldsymbol{\beta}}(Y\mid X=x,\boldsymbol{Z}){\Pr}_{\boldsymbol{\eta}}(X=x\mid X^{*},\boldsymbol{Z}){\Pr}(X^{*},\boldsymbol{Z})$ instead.
+[…]
+Queried neighborhoods are assumed to follow (1), while unqueried ones are assumed to instead follow:
+${\Pr}_{\boldsymbol{\beta},\boldsymbol{\eta}}(Y,X^{*},\boldsymbol{Z})$ $=\sum_{x=0}^{1}{\Pr}_{\boldsymbol{\beta}}(Y\mid X=x,\boldsymbol{Z}){\Pr}_{\boldsymbol{\eta}}(X=x\mid X^{*},\boldsymbol{Z}){\Pr}(X^{*},\boldsymbol{Z}).$
 2.2 Maximum Likelihood Estimation
@@
 $\propto\prod_{i=1}^{N}\{{\Pr}_{\boldsymbol{\beta}}(Y_{i}\mid X_{i},\boldsymbol{Z}_{i}){\Pr}_{\boldsymbol{\eta}}(X_{i}\mid X_{i}^{*},\boldsymbol{Z}_{i})\}^{Q_{i}}$
-×$\times\left\{\sum_{x=0}^{1}{\Pr}_{\boldsymbol{\beta}}(Y_{i}\mid X_{i}=x,\boldsymbol{Z}_{i}){\Pr}_{\boldsymbol{\eta}}(X_{i}=x\mid X_{i}^{*},\boldsymbol{Z}_{i})\right\}^{1-Q_{i}}.$ (2)
-The query indicator controls which probability function governs observation $i$ in (2); it evaluates to $1$ for queried neighborhoods (i.e., when $X_{i}$ is available) and to 00 otherwise.
+[…]
+The query indicator controls which probability function governs observation $i$ in (2); it evaluates to $1$ for queried neighborhoods (i.e., when $X_{i}$ is available) and to 0 otherwise.
 Notice that the $N$ neighborhoods are assumed to be independent in this likelihood.
 Modifying the likelihood to allow spatial autocorrelation is a promising future direction.
-The MLEs $\hat{\boldsymbol{\beta}}$ maximize (2) and are found via an EM algorithm (Web Appendix A.2).
+The MLEs $\hat{\boldsymbol{\beta}}$ maximize (2) and are found via an EM algorithm (Web Appendix A.2 and Supplemental Figure S1).
 Standard errors (SEs) for $\hat{\boldsymbol{\beta}}$ are conveniently estimated using numerical derivatives.
@@
 3 Simulations
-We conducted simulation studies to test the performance of our MLE against the gold standard, naive, and complete case analyses.
+We conducted simulation studies to test the performance of our MLEs against the gold standard, naive, and complete case analyses.
 The gold standard analysis fit a Poisson regression using $X$ for all observations, while the naive analysis used error-prone version $X^{*}$, and the complete case used $X$ for only the queried subset.
 Data generation was designed to mimic our data from the Piedmont Triad (Section 4), with primary variations of the sample size $N$, query proportion $q$, and positive predictive value $PPV=\Pr(X=1|X^{*}=1,Z)$.
-[…]
-We focus here on simulating varied misclassification rates (Section 3.1) and query percentages (Section 3.2).
+We considered samples of $N=390$ neighborhoods, approximately the number of census tracts in the Piedmont Triad, or $N=2200$ neighborhoods, approximately the number of census tracts in North Carolina.
+In all settings, we first simulated a covariate ($Z$), followed by a potentially misclassified exposure ($X^{*}\mid Z$), a true exposure ($X\mid X^{*},Z$), an offset ($O$), and a count outcome ($Y\mid X,Z$).
+Nuisance parameters were chosen to force a prespecified $PPV$.
+Based on query proportion $q$ ($q\in[0,1]$), a subset of $n=Nq$ neighborhoods had non-missing $X$ for the complete case and MLE analyses.
+All $n$ were chosen from those with $X^{*}=1$.
+Since false negatives for food access are not possible, neighborhoods with $X^{*}=0$ must have $X=0$ and thus do not need to be queried.
+For more details on simulation setup, see Web Appendix B.1.
+We ran $\mathrm{1000}$ replications per setting and recorded summary metrics capturing empirical bias, variance, and relative efficiency for $\hat{\beta}_{1}$, the adjusted log PR on $X$.
+We also assessed the empirical coverage probability (CP) of the 95% Wald confidence interval and the validity of the SE estimator.
+We focus here on results under varied misclassification rates (Section 3.1) and query percentages (Section 3.2).
 Unless otherwise stated, we fixed $q=0.1$, $PPV=0.6$, $\beta_{1}=0.18$, and outcome prevalence $\approx 10\%$.
-Additional results with two-sided misclassification in $X^{*}$ and varied outcome prevalence can be found in Web Appendix B.2–B.3.
-Like in the cases below, the MLE remained unbiased and more efficient than the complete case analysis.
+Additional results with two-sided misclassification in $X^{*}$ and varied outcome prevalence can be found in Web Appendices B.2–B.3.
+Like the settings below, the MLE remained unbiased and more efficient than the complete case analysis.
 3.1 Results Under Varied Misclassification Rates
-We considered $50\%$ to $90\%$ $PPV$ at sample sizes of $N=390$ and $2200$.
+We considered $PPV$ from $0.5$ to $0.9$ at sample sizes of $N=390$ and $2200$.
 The bias of the MLE remained comparable to that of the gold standard and complete case analyses in all settings (Table 1).
 Even under the most severe misclassification ($PPV=0.5$), the MLE was less than $0.1\%$ biased.
-However, the naive analysis yielded heavy downward bias ranging from $-8.9\%$ to $-46.5\%$, worsening as $PPV$ decreased and persisting in larger samples.
-Table 1: Simulation results for $\hat{\beta}_{1}$, the adjusted log prevalence ratio on $X$, under varied positive predictive value ($PPV$) and sample size ($N$), assuming $\approx 10\%$ outcome prevalence, $\beta_{1}=0.18$, and $q=0.1$ query percentage.
+However, the naive analysis yielded heavy downward bias ranging from $9\%$ to $47\%$, worsening as $PPV$ decreased and persisting in larger samples.
+Table 1: Simulation results for $\hat{\beta}_{1}$, the adjusted log PR on $X$, under varied positive predictive value (PPVPPV) and sample size (NN), assuming $\approx 10\%$ outcome prevalence, $\beta_{1}=0.18$, and $q=0.1$ query percentage.
 Gold Standard Naive Complete Case Maximum Likelihood Estimator
 $\boldsymbol{N}$ $\boldsymbol{PPV}$ Bias ESE Bias ESE Bias ESE RE Bias ESE ASE CP RE
-390 0.5 $0.000$ $0.005$ $-0.465$ $0.008$ $0.002$ $0.015$ $0.098$ $0.000$ $0.005$ $0.005$ $0.949$ $0.955$
-0.6 $0.000$ $0.005$ $-0.367$ $0.007$ $0.000$ $0.016$ $0.085$ $0.000$ $0.005$ $0.005$ $0.954$ $0.975$
-0.7 $0.001$ $0.005$ $-0.271$ $0.007$ $-0.001$ $0.018$ $0.075$ $0.001$ $0.005$ $0.005$ $0.938$ $0.985$
-0.8 $0.000$ $0.005$ $-0.181$ $0.007$ $0.000$ $0.019$ $0.059$ $0.000$ $0.005$ $0.005$ $0.953$ $0.985$
-0.9 $0.000$ $0.005$ $-0.089$ $0.006$ $0.000$ $0.030$ $0.028$ $0.000$ $0.005$ $0.005$ $0.944$ $0.987$
-2200 0.5 $0.000$ $0.002$ $-0.465$ $0.003$ $-0.001$ $0.006$ $0.107$ $0.000$ $0.002$ $0.002$ $0.943$ $0.986$
-0.6 $0.001$ $0.002$ $-0.366$ $0.003$ $0.001$ $0.007$ $0.097$ $0.001$ $0.002$ $0.002$ $0.935$ $0.984$
-0.7 $0.000$ $0.002$ $-0.272$ $0.003$ $-0.002$ $0.007$ $0.087$ $0.000$ $0.002$ $0.002$ $0.951$ $0.992$
-0.8 $0.000$ $0.002$ $-0.179$ $0.003$ $0.000$ $0.008$ $0.065$ $0.000$ $0.002$ $0.002$ $0.948$ $1.002$
-0.9 $-0.001$ $0.002$ $-0.089$ $0.003$ $-0.002$ $0.011$ $0.036$ $-0.001$ $0.002$ $0.002$ $0.944$ $0.992$
-[…]
+390 0.5 $0.000$ $0.005$ 0.465-0.465 $0.008$ $0.002$ $0.015$ $0.098$ $0.000$ $0.005$ $0.005$ $0.949$ $0.955$
+0.6 $0.000$ $0.005$ 0.367-0.367 $0.007$ $0.000$ $0.016$ $0.085$ $0.000$ $0.005$ $0.005$ $0.954$ $0.975$
+0.7 $0.001$ $0.005$ 0.271-0.271 $0.007$ 0.001-0.001 $0.018$ $0.075$ $0.001$ $0.005$ $0.005$ $0.938$ $0.985$
+0.8 $0.000$ $0.005$ 0.181-0.181 $0.007$ $0.000$ $0.019$ $0.059$ $0.000$ $0.005$ $0.005$ $0.953$ $0.985$
+0.9 $0.000$ $0.005$ 0.089-0.089 $0.006$ $0.000$ $0.030$ $0.028$ $0.000$ $0.005$ $0.005$ $0.944$ $0.987$
+2200 0.5 $0.000$ $0.002$ 0.465-0.465 $0.003$ 0.001-0.001 $0.006$ $0.107$ $0.000$ $0.002$ $0.002$ $0.943$ $0.986$
+0.6 $0.001$ $0.002$ 0.366-0.366 $0.003$ $0.001$ $0.007$ $0.097$ $0.001$ $0.002$ $0.002$ $0.935$ $0.984$
+0.7 $0.000$ $0.002$ 0.272-0.272 $0.003$ 0.002-0.002 $0.007$ $0.087$ $0.000$ $0.002$ $0.002$ $0.951$ $0.992$
+0.8 $0.000$ $0.002$ 0.179-0.179 $0.003$ $0.000$ $0.008$ $0.065$ $0.000$ $0.002$ $0.002$ $0.948$ $1.002$
+0.9 0.001-0.001 $0.002$ 0.089-0.089 $0.003$ 0.002-0.002 $0.011$ $0.036$ 0.001-0.001 $0.002$ $0.002$ $0.944$ $0.992$
+  • Note: Bias and ESE are, respectively, the empirical relative bias and standard error of each estimator; ASE is the average of the standard error estimator; CP is the empirical coverage probability of the 95% Wald confidence intervals (computed with the standard error estimator); and RE is the relative efficiency of the estimator to the gold standard.
+All entries are based on $\mathrm{1000}$ replicates.
+Across all settings, 22 replicates of the MLE ($0.2\%$) were replaced with the naive analysis due to $100\%$ PPV in the queried subset and perfect separation in the misclassification model.
 The empirical SEs of the MLE were smaller than those of the complete case analysis in every setting.
 In addition, the average SEs of the MLE approximated the empirical versions very well, which also led to empirical CP in the correct neighborhood of 95%.
-Finally, the relative efficiency of the MLE to the gold standard analysis was over five times higher than that of the complete case analysis in every setting (between 0.960.960.960.961.001.001.001.00 versus 0.040.040.040.040.110.110.110.11).
+Finally, the relative efficiency of the MLE to the gold standard analysis was over five times higher than that of the complete case analysis in every setting (between 0.960.961.001.00 versus 0.040.040.110.11).
 3.2 Results Under Varied Query Percentages
-We also considered querying proportions between $q=10\%$ and $75\%$ from samples $N=390$ and $2200$ neighborhoods.
+We also considered querying proportions between $q=0.1$ and $0.5$ from samples of $N=390$ and $2200$ neighborhoods.
 The bias of the MLE remained comparable to that of the gold standard analysis; both methods were under 0.5% biased in all settings (Table 2).
@@
 Again, the average SEs of the MLE matched their empirical counterparts, and empirical CP was close to the nominal 95% level.
-Table 2: Simulation results for $\beta_{1}$ under varied query percentage ($q$) and sample size ($N$), assuming $\approx 10\%$ outcome prevalence, $\beta_{1}=0.18$, and positive predictive value $=0.6$.
+Table 2: Simulation results for $\hat{\beta}_{1}$ under varied query percentage ($q$) and sample size (NN), assuming $\approx 10\%$ outcome prevalence, $\beta_{1}=0.18$, and positive predictive value =0.6=0.6.
 Gold Standard Naive Complete Case Maximum Likelihood Estimator
 $\boldsymbol{N}$ $\boldsymbol{q}$ Bias ESE Bias ESE Bias ESE RE Bias ESE ASE CP RE
-390 0.10 $0.000$ $0.005$ $-0.367$ $0.007$ $0.000$ $0.016$ $0.085$ $0.000$ $0.005$ $0.005$ $0.954$ $0.975$
-0.25 $0.002$ $0.005$ $-0.367$ $0.007$ $0.003$ $0.010$ $0.227$ $0.002$ $0.005$ $0.005$ $0.953$ $0.974$
-0.50 $0.001$ $0.005$ $-0.366$ $0.007$ $0.002$ $0.007$ $0.463$ $0.001$ $0.005$ $0.005$ $0.953$ $0.992$
-0.75 $0.000$ $0.005$ $-0.367$ $0.007$ $0.000$ $0.006$ $0.673$ $0.000$ $0.005$ $0.005$ $0.953$ $1.000$
-2200 0.10 $0.001$ $0.002$ $-0.366$ $0.003$ $0.001$ $0.007$ $0.097$ $0.001$ $0.002$ $0.002$ $0.935$ $0.984$
-0.25 $0.000$ $0.002$ $-0.368$ $0.003$ $0.000$ $0.004$ $0.227$ $0.000$ $0.002$ $0.002$ $0.949$ $0.980$
-0.50 $0.000$ $0.002$ $-0.367$ $0.003$ $0.000$ $0.003$ $0.508$ $0.000$ $0.002$ $0.002$ $0.945$ $0.992$
-0.75 $-0.001$ $0.002$ $-0.368$ $0.003$ $-0.001$ $0.002$ $0.695$ $-0.001$ $0.002$ $0.002$ $0.948$ $1.000$
-[…]
+390 0.10 $0.000$ $0.005$ 0.367-0.367 $0.007$ $0.000$ $0.016$ $0.085$ $0.000$ $0.005$ $0.005$ $0.954$ $0.975$
+0.25 $0.002$ $0.005$ 0.367-0.367 $0.007$ $0.003$ $0.010$ $0.227$ $0.002$ $0.005$ $0.005$ $0.953$ $0.974$
+0.50 $0.001$ $0.005$ 0.366-0.366 $0.007$ $0.002$ $0.007$ $0.463$ $0.001$ $0.005$ $0.005$ $0.953$ $0.992$
+2200 0.10 $0.001$ $0.002$ 0.366-0.366 $0.003$ $0.001$ $0.007$ $0.097$ $0.001$ $0.002$ $0.002$ $0.935$ $0.984$
+0.25 $0.000$ $0.002$ 0.368-0.368 $0.003$ $0.000$ $0.004$ $0.227$ $0.000$ $0.002$ $0.002$ $0.949$ $0.980$
+0.50 $0.000$ $0.002$ 0.367-0.367 $0.003$ $0.000$ $0.003$ $0.508$ $0.000$ $0.002$ $0.002$ $0.945$ $0.992$
+  • Note: Bias and ESE are, respectively, the empirical relative bias and standard error of each estimator; ASE is the average of the standard error estimator; CP is the empirical coverage probability of the 95% Wald confidence intervals (computed with the standard error estimator); and RE is the relative efficiency of the estimator as compared to the gold standard.
+All entries are based on $\mathrm{1000}$ replicates.
 The relative efficiency of the MLE to the gold standard analysis exceeded that of the complete case analysis in every setting.
@@
 Using one-mile and half-mile radii, we discretized these distances to create our exposures $X$ and $X^{*}$.
-For example, $X=1$ if a census tract has at least one healthy food retailer within the threshold and $=0$ otherwise.
+For example, $X=1$ if a census tract had at least one healthy food retailer within the threshold and =0=0 otherwise.
 Queried data (i.e., with both $X$ and $X^{*}$) were available for all census tracts.
-However, to demonstrate the two-phase design, we assumed that only a subset had available route-based access data for the MLE and complete case analyses.
+However, to demonstrate the two-phase design, we assumed that only a subset had available route-based access data for the MLEs and complete case analyses.
 After running the gold standard analysis on the full data, we created a partially queried version wherein $X$ was only available for census tracts with $X^{*}=0$ (such that $X=0$ necessarily) and a subset of those with $X^{*}=1$.
-A random sample of $n=77$ tracts with $X^{*}=1$, targeting a 50/50 split between metropolitan/non-metropolitan, was chosen.
-Only tracts with error-prone access (i.e., $X^{*}=1$) needed to be validated, since those with $X^{*}=0$ were already known to have $X=0$ and do not inform the misclassification mechanism.
+A random sample of $n=77$ tracts with $X^{*}=1$, targeting a 50/50 split between metropolitan/non-metropolitan, was chosen (Supplemental Figure S2).
+Only tracts with error-prone access (i.e., $X^{*}=1$) needed to be validated, since those with $X^{*}=0$ must have $X=0$ and do not inform the misclassification mechanism.
 The naive analysis used the error-prone access variable on the entire sample, and the complete case analysis used error-free access variable on the subset of queried data.
-The MLE used the error-prone and error-free access indicators when available and just the error-prone version otherwise.
+The MLEs used the error-prone and error-free access indicators when available and just the error-prone version otherwise.
 4.2 Diabetes Prevalence and Healthy Food Access
-At both half- and one-mile radii, the majority of tracts in the Piedmont Triad ($n=353$ [$91\%$] and $n=265$ [$68\%$], respectively) did not have access to healthy foods (Figure 1).
+At both half- and one-mile radii, most tracts ($n=353$ $91\%$] and $n=265$ $68\%$], respectively) did not have access to healthy foods (Figure 1).
 Larger, rural tracts tended not to have access.
 Smaller, urban tracts tended to have access, or at least to be misclassified as such based on Haversine distances.
-Of the $192$ tracts with access to healthy foods based on one-mile Haversine distances, only $122$ actually had it based on route-based ones ($PPV=122/192=64\%$).
-As expected, there was more misclassification at the smaller threshold ($PPV=34/85=40\%$).
+Of the $192$ tracts with access to healthy foods based on one-mile Haversine distances, only $122$ actually had it based on route-based ones ($PPV=64\%$).
+As expected, the smaller threshold saw more misclassification ($PPV=40\%$).
 Figure 1: Map of Piedmont Triad census tracts’ access at one-half and one mile thresholds.
-The estimated prevalence ratio $\widehat{PR}_{non}=\exp(\hat{\beta_{1}})$ compared the predicted prevalence of diabetes in non-metropolitan census tracts with versus without access to healthy foods.
-In metropolitan census tracts, $\widehat{PR}_{metro}=\exp(\hat{\beta_{1}}+\hat{\beta_{3}})$ was interpreted similarly, and 95% CIs for this prevalence ratio were calculated as $\exp\left\{(\hat{\beta_{1}}+\hat{\beta_{3}})\pm 1.96\widehat{\textrm{SE}}(\hat{\beta_{1}}+\hat{\beta_{3}})\right\}$, where $\widehat{\textrm{SE}}(\hat{\beta_{1}}+\hat{\beta_{3}})=\sqrt{\widehat{\textrm{V}}(\hat{\beta_{1}})+\widehat{\textrm{V}}(\hat{\beta_{3}})+\widehat{\textrm{Cov}}(\hat{\beta_{1}},\hat{\beta_{3}})}$.
-We focus on these two prevalence ratios; the full models are summarized in Supplemental Table S4.
-At either threshold, prevalence of diabetes in neighborhoods with access to healthy foods was expected to be higher than in those without (Figure 2).
-This counterintuitive trend held at both distance thresholds and across all analysis methods.
-For example, based on the gold standard, diabetes prevalence in metropolitan tracts with access to healthy foods at one mile was expected to be 18% higher than in those without ($\widehat{PR}_{metro}=1.18$; 95% CI: $1.16$, $1.21$); among non-metropolitan tracts, those with access were expected to have 9% higher prevalence ($\widehat{PR}_{non}=1.09$; 95% CI: $1.06$, $1.11$).
+[…]
+At either threshold, diabetes prevalence in neighborhoods with access to healthy foods was expected to be higher than in those without (Figure 2).
+This trend seems somewhat counterintuitive, but the literature does not agree on a clear pattern linking food access and diabetes (e.g., Gucciardi et al., 2014 versus Flint et al., 2020).
+Potential confounding (e.g., by socioeconomic status) or uncertainty around the outcome data (small area estimates) may play a role in this debate.
 Figure 2: Prevalence ratio estimates and 95% confidence intervals for healthy food access at a one-half and one mile threshold in the Piedmont Triad using four analysis methods.
+The observed effect direction held at both distance thresholds and across all analysis methods.
+For example, based on the gold standard, diabetes prevalence in metropolitan tracts with access to healthy foods at one mile was expected to be 18% higher than in those without ($\widehat{\textrm{PR}}_{metro}=1.18$; 95% CI: $1.16$, $1.21$); among non-metropolitan tracts, those with access were expected to have 9% higher prevalence ($\widehat{\textrm{PR}}_{non}=1.09$; 95% CI: $1.06$, $1.11$).
 The difference between tracts with and without access was more pronounced in metropolitan areas, according to the gold standard analyses.
 At both thresholds, the MLE also captured this difference, while the complete case analysis only did so at the one-half mile radius, and the naive analysis did not catch it at either threshold.
-The MLE actually exaggerated this difference at the one-mile threshold but otherwise closely aligned with the gold standard.
+The MLEs actually exaggerated this difference at the one-mile threshold but otherwise closely aligned with the gold standard analysis.
 Its 95% confidence intervals were slightly narrower than the complete case.
-For more details, see Web Appendix C.
+For more details and further sensitivity analysis, see Web Appendix C.
 5 Conclusion
 We introduced a novel maximum likelihood estimator for Poisson regression with a misclassified binary exposure.
-In simulations, we observe that the MLE, on average, corrects for bias induced by the misclassified exposure in the naive analysis and reduces variance induced by deleting unqueried data in the complete case analysis.
-In our analysis of the Piedmont Triad, tracts with access to healthy foods had higher diabetes prevalence, on average, than those without.
+In simulations, the MLE, on average, corrected for bias induced by the misclassified exposure in the naive analysis and reduced variance induced by deleting unqueried data in the complete case analysis.
+In our analysis of the Piedmont Triad, census tracts with access to healthy foods had higher diabetes prevalence, on average, than those without.
 This difference was more pronounced for metropolitan tracts.
@@
 We are potentially limited by our choice to assume a parametric logistic regression model for the exposure error mechanism.
-However, this choice has been shown to be fairly robust to misspecification (Lotspeich et al., 2022).
+However, this model has been shown to be fairly robust to misspecification (Lotspeich et al., 2022).
 Still, nonparametric or semiparametric error models are promising.
@@
 Census tracts do not always fairly describe neighborhoods, but they were the most granular level of available diabetes data.
-Additionally, the model used to generate the small area diabetes estimates incorporates many socioeconomic factors that may have been relevant to our analysis but could not be included because of “double dipping.”
-There are many interesting avenues of further exploration.
-It may be informative to explore spatial relationships among adjacent tracts, vary the parametric form of the outcome model, or modify the MLE to handle other types of data.
+Other studies (e.g., Rummo et al., 2021) consider food access at the individual level.
+If we had individual-level data, our proposed methods would still apply, simply omitting the model offset.
+Additionally, the model used to generate the small area diabetes estimates incorporates many socioeconomic factors that may have been relevant to our analysis but could not be included due to “double dipping.”
+In the future, it may be informative to explore spatial relationships between adjacent tracts, vary the outcome model, or modify the MLE to handle other data types.
 The MLE’s efficiency could be further improved by a more strategic query design.
-Targeted strategies for selecting the most informative observations are common in the two-phase design literature.
+Targeted strategies to select the most informative observations are common in the two-phase design literature.
+(Flint et al., 2020)
 SUPPLEMENTARY MATERIAL
@@
 Data:
-R package:
-R package possum implementing the MLE (GNU zipped tar file)
 R code:
-R scripts for the simulation studies and Piedmont Triad analysis (.zip file)
+R package possum implementing the MLE (GNU zipped tar file) and R scripts for the simulation studies and Piedmont Triad analysis (.zip file)
 Additional materials: